# BeamLLM: Vision-Empowered mmWave Beam Prediction with Large Language Models

### v1

@@ III-C Learning Phase
The beam prediction task is essentially a classification problem; therefore, the model parameters are optimized by minimizing the cross-entropy, which is expressed as:
$\mathcal{L}=\sum_{j=t}^{t+H-1}\sum_{m=1}^{|\mathbf{\mathcal{F}}|}f^{*}[j]\log_{2}(p_{m}[j]),$ (9)
where $f^{*}_{m}[j]\in\{0,1\}^{M}$ is the $m$-th element of the one-hot encoded vector of $\mathbf{f}^{*}_{m}[j]$ and $p_{m}[j]$ is the $m$-th element of the output vector $\hat{\mathbf{p}}[j]$ at time step $j$, respectively.
@@ IV-A Experimental Settings
Dataset Processing: We adopt Scenario 8 of the DeepSense 6G dataset for our simulation, which simulates a V2I mmWave communication setup. The BS is equipped with an RGB camera and a $16$-element $60$ GHz mmWave phased array, while the mobile UE serves as a mmWave transmitter. During data collection, the UE passes by the BS multiple times. At each time step, the BS captures an RGB image of the UE while scanning all predefined beams and measuring the received power for all $|\mathcal{F}|=32$ beams in a codebook. The multimodal data streams are synchronized to ensure temporal consistency.
The dataset is split into $70\%$ training, $10\%$ validation, and $20\%$ test sets. The dataset consists of multiple data sequences. In each data sequence, the vehicle passes by the BS once. Each data sequence is a pair comprising an RGB image sequence and a beam index sequence. For each data sequence, we decompose it into data samples using a sliding window of size $13$. As previously mentioned, during training, we use an observation window of size $T$, and we train the model to predict future beams over a horizon $H$. Therefore, the input to the encoder for the model is ${\mathbf{X}_{I}[1],\ldots,\mathbf{X}_{I}[T]}$. In both beam prediction methods, the expected output from the decoder is ${\hat{\mathbf{p}}[T+1],\ldots,\hat{\mathbf{p}}[T+H]}$. Since we maintain a fixed sequence length of 13, we set $T=8,H=5$ as standard prediction and $T=3,H=10$ as few-shot prediction.
Baselines: We compare our approach with several classical time-series models, including RNN[1], GRU, and LSTM. Additionally, to validate the effectiveness of the PaP module, we conduct an ablation study by comparing our model with and without PaP in the standard prediction setup.
Parameter Settings: BeamLLM is configured as following: 1) A widely-used language model, i.e., GPT-$2$[16], is employed as the LLM backbone; 2) It is trained with Adam optimizer, where the batch size and initial learning rate (LR) are $16$ and $0.001$, respectively. Additionally, a multi-step LR scheduler in $1$, $5$, $10$, $15$, $20$, $25$, $30$, $40$ epochs with a decay factor of $\gamma=0.9$ is employed; 3) The training process is set to $200$ epochs. The detailed model parameters are shown in Table I.
TABLE I: Parameter Settings of Different Models
LLM RNN, GRU, LSTM
Patch Reprogramming Output Projection Embedding Layer Sequence Model
Same as [12] except $V^{\prime}=64$
Linear 1: $4\times 8$
ReLU
Linear 2: $8\times 16$
ReLU
Linear 3: $16\times 32$
Softmax
Linear: $4\times 32$
Layer $1-3$: $32\times 32$
Linear: $32\times 32$
@@
where $N_{S}$ represents the total number of samples in the test set, $m_{i}$ denotes the index of the ground truth optimal beam for the $i$-th sample, and $Q_{k}$ is the set of indices for the top-$K$ predicted beams, sorted by the element values in $\hat{\mathbf{P}}$ for each time sample.
IV-B Standard Prediction
In Figs. 3 and 4, we present a comparative analysis of the top-$1$ and top-$3$ accuracy in the standard predictions across all models. Increasing $K$ improves top-$K$ accuracy, while as prediction horizon extends further into the future, the accuracy gradually decreases. Among the models, BeamLLM achieves the highest top-$1$ and top-$3$ accuracy scores, reaching $\mathbf{61.01\%}$ and $\mathbf{97.39\%}$, respectively. Additionally, as the number of time samples increases, the decay in the top-$K$ accuracy for the LSTM model is minimal. Specifically, the top-$1$ and top-$3$ accuracy only decrease by $\mathbf{6.03\%}$ and $\mathbf{1.65\%}$, respectively, across time samples ranging from $1$ to $5$. This smaller reduction highlights the adaptability of LSTMs, as their gating mechanism adjusts information retention and updating based on task demands.
@@ IV-B Standard Prediction
Figure 3: Top-$1$ accuracy performance of the proposed method comparing to several baselines in the standard prediction task.
Figure 4: Top-$3$ accuracy performance of the proposed method comparing to several baselines in the standard prediction task.
IV-C Few-Shot Prediction
In Figs. 5 and 6, we present the top-$1$ and top-$3$ accuracy performance for the few-shot forecasting task. Existing DL prediction methods perform poorly in this scenario, particularly as the prediction horizon extends, resulting in severe performance degradation. Even for the previously most stable LSTM model, during the progression from time sample $1$ to 10, the top-$1$ accuracy is decreased by $\mathbf{16.48\%}$, and the top-$3$ accuracy is decreased by $\mathbf{11.58\%}$. In contrast, BeamLLM significantly outperforms all baseline methods, with only $\mathbf{12.56\%}$ and $\mathbf{5.55\%}$ performance degradation, respectively. We attribute this superior performance to the successful activation of knowledge through the reprogrammed LLM.
Figure 5: Top-$1$ accuracy performance of the proposed method comparing to several baselines in the few-shot prediction task.
Figure 6: Top-$3$ accuracy performance of the proposed method comparing to several baselines in the few-shot prediction task.
IV-D Analysis of Reprogramming
We provide a case study of reprogramming 64 time series patches with 64 text prototypes, as shown in Fig. 7. The figure consists of three subplots, each visualizing the similarity between text prototypes computed as the scaled dot product ${\mathbf{Q}_{k}^{(i)}{\mathbf{K}_{k}^{(i)}}^{\top}}/{\sqrt{d_{k}}}$, across distinct training epochs. A color bar accompanies three subplots, with values ranging from 0 (dark purple, denoting low similarity) to 1 (bright yellow, denoting high similarity). The observed transition from a noisy, scattered pattern at epoch $1$ to a sparse and concentrated representation by epoch $5$ illustrates that the learned prototypes effectively capture the local semantic information of the input features. Moreover, the heatmaps indicate that only a few text prototypes exhibit significant correlations with the patches, suggesting BeamLLM’s ability to adaptively prioritize prototypes most relevant to the local semantic context.
Figure 7: A showcase of text prototype evolution across training epochs.
IV-E Analysis of Complexity
All experiments are conducted in the same environment, specifically on Google Colab with an NVIDIA A100100100100 GPU and $40$ GB of RAM. We investigate the training complexity and inference complexity of different models in terms of the number of trainable and non-trainable parameters, as well as the average inference time per epoch. From Table II, we observe that although the backbone model is frozen, the number of trainable parameters of BeamLLM remains large. Meanwhile, the average inference time is significantly longer than that of traditional models. While its high deployment cost poses a challenge, this also indicates that the full potential of BeamLLM has yet to be fully explored.
TABLE II: The Number of Model Parameters and Average Inference Time
Models # of trainable parameters # of non-trainable parameters Average inf. time (sec)
RNN $8,641$ 00 $0.17$
GRU $18,641$ 00 $0.15$
LSTM $26,593$ 00 $0.25$
BeamLLM $130,056,118$ $124,439,808$ $10.85$
@@ V Conclusions
However, the performance gains come with increased inference complexity. The massive parameter scale of LLMs may introduce higher resource consumption and latency. Nevertheless, BeamLLM remains practical, particularly due to its exceptional few-shot prediction capability, which enables predictions over a longer horizon. Practical deployments require a trade-off between model complexity and real-time constraints, necessitating optimizations such as model compression or lightweight architecture design. By advancing these aspects, the proposed framework could serve as a scalable and efficient beam management solution for 6G ISAC systems.

### v2

@@ III-C Learning Phase
As shown in Fig. 2, certain modules are marked with a flame icon, indicating that they are trainable components. In contrast, the LLM backbone is marked with a snowflake icon, signifying that its parameters are frozen during training. The model is trained in a supervised manner using image inputs paired with the corresponding ground truth optimal beam indices. The beam prediction task is essentially a time-series forecasting task, where each time step corresponds to a classification task; therefore, the model parameters are optimized by minimizing the cross-entropy, which is expressed as:
$\mathcal{L}=-\sum_{j=t}^{t+T_{\text{pred}}-1}\sum_{m=1}^{M}\mathbbm{1}\{\hat{m}^{*}[j]=m^{*}[j]\}\log_{2}(\hat{p}_{m}[j]),$ (9)
where $\hat{p}_{m}[j]$ is the $m$-th element of the output vector $\hat{\mathbf{p}}[j]$ at time step $j$, respectively.
@@ IV-A Experimental Settings
Dataset Processing: We adopt Scenario $8$ of the DeepSense $6$G dataset for our simulation, which simulates a V2222I mmWave communication setup. The BS is equipped with an RGB camera and a $16$-element $60$ GHz mmWave phased array, while the mobile UE serves as a mmWave transmitter. During data collection, the UE passes by the BS multiple times. At each time step, the BS captures an RGB image of the UE while scanning all predefined beams and measuring the received power for all $M=32$ beams in a codebook. The multimodal data streams are synchronized to ensure temporal consistency.
The dataset is split into $70\%$ training, $10\%$ validation, and $20\%$ test sets. The dataset consists of multiple data sequences. Each data sequence is a pair comprising an RGB image sequence and a beam index sequence. For each data sequence, we decompose it into data samples using a sliding window of size $13$. As previously mentioned, during training, we use an observation window of size $T_{\text{hist}}$, and we train the model to predict future beams over a horizon $T_{\text{pred}}$. Therefore, the input to the encoder for the model is $\mathbf{X}_{I}[t-T_{\text{hist}}],\ldots,\mathbf{X}_{I}[t-1]$. In both beam prediction methods, the expected output from the decoder is $\hat{\mathbf{p}}[t],\ldots,\hat{\mathbf{p}}[t+T_{\text{pred}}-1]$. Since we maintain a fixed sequence length of $13$, we set $T_{\text{hist}}=8$, $T_{\text{pred}}=5$ as standard prediction and $T_{\text{hist}}=3$, $T_{\text{pred}}=10$ as few-shot prediction.
Baselines: We compare our approach with several classical time-series models, including RNN (the implementation is consistent with [1]), GRU, and LSTM. Additionally, to validate the effectiveness of the PaP module, we conduct an ablation study by comparing our model with and without PaP in the standard prediction setup.
Parameter Settings: BeamLLM is configured as following: 1) A widely-used language model, i.e., GPT-$2$[15], is employed as the LLM backbone; 2) It is trained with Adam optimizer, where the batch size and initial learning rate (LR) are $16$ and $0.01$, respectively. Additionally, a multi-step LR scheduler in $1$, $5$, $10$, $15$, $20$, $25$, $30$, $40$ epochs with a decay factor of $\gamma=0.9$ is employed; 3) The training process is set to $200$ epochs. The detailed model parameters are shown in Table I.
TABLE I: Parameter settings of different models.
Component BeamLLM RNN/GRU/LSTM
Input Layers
Patch Reprogramming
(Same as [11], $V^{\prime}=64$)
Linear $4\times 32$
Output Layers
Output Projection:
1. Linear $4\times 16$ + ReLU
2. Linear $16\times 32$ + ReLU
3. Linear $32\times 32$
Sequence Model:
Layers $1$-$4$: $32\times 32$
@@ 3. Linear $32\times 32$
where $N_{\text{test}}$ represents the total number of samples in the test set, $m_{i}$ denotes the index of the ground truth optimal beam for the $i$-th sample, and $Q_{k}$ is the set of indices for the top-$K$ predicted beams, sorted by the element values in $\hat{\mathbf{P}}$ for each time sample.
IV-B Dataset Visualization
Figure 3: Comparison of power profiles for the first $3$ samples associated with optimal beams $20$, $40$, and $60$, respectively.
(a) Vision data visualization for the first $3$ samples associated with optimal beam $20$.
(b) Vision data visualization for the first $3$ samples associated with optimal beam $40$.
(c) Vision data visualization for the first $3$ samples associated with optimal beam $60$.
Figure 4: Vision data visualization for the first $3$ samples associated with optimal beams $20$, $40$, and $60$, respectively.
Fig. 3 illustrates the relative received power distribution across various beam indices ($20$, $40$, $60$) for the first three samples, highlighting how peak power varies with different beam selections. Note that due to the complexity of absolute power calibration, the dataset provides dimensionless relative received signal power [14]. From the figure, it can be seen that as the optimal beam index increases, the relative received power shows an overall increasing trend. This implies that the distance between transceivers in the dataset may be smaller when the beam index is larger, leading to an increase in received power. Fig. 4 shows the first three image samples when the optimal beam is $20$, $40$, and $60$, respectively. We can see that as the blue car travels from the left side of the screen to the right, the optimal beam index progressively increases. This suggests that the optimal beam is highly dependent on the direction and position of the UE, and that visual information can intuitively provide information on the relative orientation position of the UE, as well as additional environmental information.
IV-C Convergence Analysis
Fig. 5 presents the training dynamics of BeamLLM over $200$ epochs, where the left vertical axis denotes the training and validation loss, and the right vertical axis represents the training accuracy. Note that training accuracy refers to the average top-$1$ accuracy in the training set. The convergence of both training and validation losses around similar levels suggests that the model has not severely overfitted. The gap between training and validation loss is not large, indicating that the model maintains a certain level of robustness on unseen data.
Figure 5: The curves of the training loss, validation loss, and training accuracy for BeamLLM.
IV-D Standard Prediction
Figure 6: Top-$K$ accuracy performance of the proposed method comparing to several baselines in the standard prediction task.
Figure 7: Top-$K$ accuracy performance of the proposed method comparing to several baselines in the few-shot prediction task.
In Fig. 6, we present a comparative analysis of the top-$1$ and top-$3$ accuracy in the standard predictions across all models. Increasing $K$ improves top-$K$ accuracy, while as prediction horizon extends further into the future, the accuracy gradually decreases. Among the models, BeamLLM achieves the highest top-$1$ and top-$3$ accuracy scores, reaching $\mathbf{61.01\%}$ and $\mathbf{97.39\%}$, respectively. Additionally, as the number of time samples increases, the decay in the top-$K$ accuracy for the LSTM model is minimal. Specifically, the top-$1$ and top-$3$ accuracy only decrease by $\mathbf{6.03\%}$ and $\mathbf{1.65\%}$, respectively, across time samples ranging from $1$ to $5$. This smaller reduction highlights the adaptability of LSTMs, as their gating mechanism adjusts information retention and updating based on task demands.
@@ IV-D Standard Prediction
IV-E Few-Shot Prediction
In Fig. 7, we present the top-$1$ and top-$3$ accuracy performance for the few-shot forecasting task. Existing DL prediction methods perform poorly in this scenario, particularly as the prediction horizon extends, resulting in severe performance degradation. Even for the previously most stable LSTM model, during the progression from time sample $1$ to 10, the top-$1$ accuracy is decreased by $\mathbf{16.48\%}$, and the top-$3$ accuracy is decreased by $\mathbf{11.58\%}$. In contrast, BeamLLM significantly outperforms all baseline methods, with only $\mathbf{12.56\%}$ and $\mathbf{5.55\%}$ performance degradation, respectively. We attribute this superior performance to the successful activation of knowledge through the reprogrammed LLM.
IV-F Analysis of Complexity
All experiments are conducted in the same environment, specifically on Google Colab with an NVIDIA A100100100100 GPU and $40$ GB of RAM. We investigate the training complexity and inference complexity of different models in terms of the number of trainable and non-trainable parameters, as well as the average inference time per sample (without YOLOv4444 feature extraction module). Overall, the inference time of all models is significantly shorter than the sampling interval, which corresponds to $7.79$ frames per second (FPS) [14]. As shown in Table II, although the backbone model in BeamLLM is frozen during training, the number of trainable parameters remains substantial. Additionally, its average inference time is higher than that of traditional models. While this introduces a higher deployment cost, it also suggests that the full potential of BeamLLM has yet to be fully realized and optimized.
@@ V Conclusions
However, the performance gains come with increased inference complexity. The massive parameter scale of LLMs may introduce higher resource consumption and latency. Nevertheless, BeamLLM remains practical, particularly due to its exceptional few-shot prediction capability, which enables predictions over a longer horizon. Practical deployments require a trade-off between model complexity and real-time constraints, necessitating acceleration methods such as model compression or lightweight architecture design. By advancing these aspects, the proposed framework could serve as a scalable and efficient beam management solution for $6$G ISAC systems.
TABLE II: The number of model parameters and average inference time.
Models # total params. # trainable params. Average inf. time (sec)
RNN $29,505$ $29,505$ $8.0\times 10^{-6}$
GRU $79,425$ $79,425$ $3.7\times 10^{-5}$
LSTM $104,385$ $104,385$ $1.9\times 10^{-5}$
BeamLLM $178,303,798$ $53,863,990$ $7.5\times 10^{-4}$